n this experiment you will react calcium chloride and sodium carbonate in a double replacement reaction, making a solid precipitate you can filter off.

CaCl2(aq) + Na2CO3(aq) --> CaCO3(s) +2 NaCl(aq)





What is the molar mass of: (0.5 pt. ea.) (whole numbers please)

CaCl2? _______________
CaCO3? ­­­­­­­­­­­­­­­­­­­­­­­­­­______________
1
Expert's answer
2020-04-18T07:08:51-0400

Solution:

1) CaCl2 - calcium chloride

M(CaCl2) = Ar(Ca) + 2*Ar(Cl) = 40 + 2*35.5 = 111 (g/mol).

M(CaCl2) = 111 g/mol.


2) CaCO3 - calcium carbonate

M(CaCO3) = Ar(Ca) + Ar(C) + 3*Ar(O) = 40 + 12 + 3*16 = 100 (g/mol).

M(CaCO3) = 100 g/mol.


Answer:

Molar mass of CaCl2 is 111 g/mol.

Molar mass of CaCO3 is 100 g/mol.